A bag of M&MS was randomly selected from the grocery store shelf, and the color counts were: Brown 20, Red 5, Yellow 10, Orange 4, Green 5, Blue 6. Mars Inc. claims that they produce M&Ms with the following distributions Brown 30% Red 20% Yellow 20% Orange 10% Green 10% Blue 10%.

1. n = 50, is there evidence of difference between the random bag that was picked and Mars Inc.'s claims?

2. Graph the results.
Answer #1

Categories Observed Expected Brown 20 50*0.3=15 Red 5 50*0.2=10 Yellow 10 50*0.2=10 (fo-fe)2/fe (20-15)2/15 = 1.667 (5-10)2/1
(3) Test Statistics The Chi-Squared statistic is computed as follows: n X - į (O; – E;) E; = 1.667 + 2.5 + 0 + 0.2 + 0 + 0.2
Therefore the claim by Mars Inc is accepted at level of significance 0.05.

p-value = P(X> 4.567), where X~chi-square with d.f =5

therefore p-value = 0.47096